# An algebraic model for rational excisive functors

### v1

@@ 4. Excisive functors
A functor $F:{{\mathsf{Sp}}}\to{{\mathsf{Sp}}}$ is $d$-excisive if it sends strongly cocartesian $(d+1)$-cubes to cartesian cubes. One should view this as a weak form of excision, i.e., a $1$-excisive functor sends pushout diagrams to pullback diagrams. We will say that $F$ is reduced if $F(\ast)\simeq\ast$. Denote by $\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})$ the $\infty$-category of reduced $d$-excisive endofunctors of spectra [^2] For technical reasons, we must restrict the domain to the full subcategory of finite objects or equivalently, colimit preserving functors, but we shall not belabour this point, nor include it in our notation. We will also only consider reduced functors, hence drop the reduced terminology and refer to the objects of $\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})$ as $d$-excisive functors. The reduced assumption is not limiting since any functor $F:{{\mathsf{Sp}}}\to{{\mathsf{Sp}}}$ is “reduced up to a constant”, i.e., decomposes as $F(X)\simeq\overline{F}(X)\times F(\ast)$, where $\overline{F}$ is reduced..
To any $F:{{\mathsf{Sp}}}\to{{\mathsf{Sp}}}$, Goodwillie calculus assigns a sequence of approximations
[…]
in which $P_{d}F$ is the universal $d$-excisive approximation to $F$. A comprehensive account of the homotopy theory of excisive functors of spectra is given in [ABHS24, §2]. The key technical input about the $\infty$-category of $d$-excisive functors is the following.
Lemma 4.1 ([ABHS24, Theorem 2.38]).
The $\infty$-category of $d$-excisive functors is a presentably symmetric monoidal stable $\infty$-category with monoidal structure given by Day convolution and monoidal unit $P_{d}h_{\mathbb{S}}=P_{d}\Sigma^{\infty}\operatorname{\mathsf{Map}}(\mathbb{S},-)\simeq P_{d}\Sigma^{\infty}\Omega^{\infty}$.
@@
The $\infty$-category of rational $d$-excisive functors $\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}$ is a presentably symmetric monoidal stable $\infty$-category with monoidal structure given by Day convolution and monoidal unit $L_{\mathbb{Q}}P_{d}h_{\mathbb{S}}$.
A functor $F$ is $d$-homogeneous if it is $d$-excisive and $P_{d-1}F\simeq\ast$. We will denote the $\infty$-category of $d$-homogeneous functors by $\mathsf{Homog}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})$. Since the $\infty$-category of $d$-excisive functors is closed under fibres, and a $(d-1)$-excisive functor is $d$-excisive, the $d$-th layer of the tower,
@@ 5. The rational Goodwillie-Burnside ring
Our attention now turns to the role of the abstract Burnside ring $A_{\mathbb{Q}}(d)$ in Goodwillie calculus, a comprehensive account of which is given in [ABHS24, §8-9]. We work rationally, as this allows us to simplify and streamline certain arguments, but integral versions of many of the statements in this section may be found in loc. cit..
@@
$\partial_{\ast}:\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})\longrightarrow{\mathsf{SymSeq}},$
@@
$\partial_{\ast}:\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})}(P_{d}h_{\mathbb{S}},P_{d}h_{\mathbb{S}})\longrightarrow\prod_{1\leq i\leq d}\operatorname{\mathsf{Hom}}_{{{\mathsf{Sp}}}}(\partial_{i}(P_{d}h_{\mathbb{S}}),\partial_{i}(P_{d}h_{\mathbb{S}})),$
@@
$\theta_{d}:\pi_{0}{\mathsf{End}}_{\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})}(P_{d}h_{\mathbb{S}})\otimes\mathbb{Q}\longrightarrow\prod_{1\leq i\leq d}\pi_{0}\operatorname{\mathsf{Hom}}_{{{\mathsf{Sp}}}}(\partial_{i}(P_{d}h_{\mathbb{S}}),\partial_{i}(P_{d}h_{\mathbb{S}}))\otimes\mathbb{Q}\cong\mathbb{Q}^{d}$
@@
between the endomorphism ring of the monoidal unit in $\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}$ and the rational Goodwillie-Burnside ring.
@@
in which the right-most map is an isomorphism by 2.4, hence it suffices to show that $\theta_{d}:R_{\mathbb{Q}}(d)\to\mathbb{Q}^{d}$ is an isomorphism.
By an elementary degree computation [ABHS24, Lemma 9.20], there are elements $[\lambda_{m}]\in R_{\mathbb{Q}}(d)$, $1\leq m\leq d$, such that the $i$th component of $\theta_{d}[\lambda_{m}]$ is $|{\mathsf{surj}}(i,m)|$ (see 3.5). A routine linear algebra exercise shows that the vectors $\theta_{d}[\lambda_{1}],\dots,\theta_{d}[\lambda_{d}]$ are linearly independent, hence $\theta_{d}$ is surjective for all $d$.
@@
$\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{1}({{\mathsf{Sp}}},{{\mathsf{Sp}}})}(P_{1}h_{\mathbb{S}},P_{1}h_{\mathbb{S}})\xrightarrow{\partial_{1}}\operatorname{\mathsf{Hom}}_{{{\mathsf{Sp}}}}(\partial_{1}(P_{1}h_{\mathbb{S}}),\partial_{1}(P_{1}h_{\mathbb{S}})),\simeq\operatorname{\mathsf{Hom}}_{{{\mathsf{Sp}}}}(\mathbb{S},\mathbb{S}),$
@@
For the induction step, assume that $\theta_{d-1}$ is an isomorphism. Since $\partial_{i}P_{d-1}\simeq\partial_{i}$ for $11\leq i\leq d-$ 1 and $\partial_{d}P_{d-1}\simeq 0$, there is a commutative diagram
[…]
where $(i_{d-1})^{\ast}:\mathbb{Q}^{d}\to\mathbb{Q}^{d-1}$ is projection onto the first $(d-1)$ factors. The above square induces a map
@@
The element $[\lambda_{d}]$ is in the kernel of $\pi_{0}(P_{d-1})\otimes\mathbb{Q}$, hence $(\theta_{d-1}\circ\pi_{0}(P_{d-1})\otimes\mathbb{Q})[\lambda_{d}]=0$, and so, by commutativity of the above square, $\theta_{d}[\lambda_{d}]$ is in the kernel of $(i_{d-1})^{\ast}$. Since $\theta_{d}[\lambda_{d}]=(0,\dots,0,d!)$ the map $\bar{\theta}_{d}$ is non-zero and it is therefore surjective as the projection onto the first $(d-1)$ factors, $(i_{d-1})^{\ast}:\mathbb{Q}^{d}\to\mathbb{Q}^{d-1}$ has $\ker((i_{d-1})^{\ast})\cong\mathbb{Q}$.
@@
[…]
in which each horizontal map is an isomorphism and the right-hand map is discussed in 3.6. In particular, the kernel of $\pi_{0}(P_{d-1})\otimes\mathbb{Q}$ is generated by the top idempotent in the rational Goodwillie-Burnside ring.
@@
We have already seen that the left square commutes, and the right square commutes, by definition of the map on the level of abstract Burnside rings. The claim about the rational kernel follows from combining 2.8 with 5.1 to see that the induced maps on the level of vertical kernels in the above diagram are all isomorphisms and the observation of 3.5 that the idempotent $e_{d}$ is a rational multiple of $[d]$. ∎
@@ 6. Splitting rational excisive functors
The Goodwillie-Burnside ring $A_{\mathbb{Q}}({\mathsf{Epi}}_{\leq d})$ associated to the category ${\mathsf{Epi}}_{\leq d}$ is equivalent to the endomorphism ring of the unit (5.1), so the splitting of the Goodwillie-Burnside ring (2.4) will induce the required splitting of the category of rational $d$-excisive functors (6.4).
The idempotents in the Goodwillie-Burnside ring lift to idempotents in the $\infty$-category of $d$-excisive functors by applying [Lur17, Lemma 1.2.4.6, Warning 1.2.4.8] and observing that the idempotents in the Goodwillie-Burnside ring are also idempotents in the homotopy category of $d$-excisive functors (5.2).
@@
by localizing with respect to the action of the idempotent $e$ and the orthogonal idempotent $1-e$. Since the $\infty$-category of $d$-excisive functors is equivalent to modules over the monoidal unit [ABHS24, Proposition 2.35], the orthogonal idempotent splitting corresponding to $e_{d}$ may be equivalently phrased as
$\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq{\mathsf{Mod}}_{\mathbb{Q}}(P_{d}h_{\mathbb{S}})\simeq{\mathsf{Mod}}_{\mathbb{Q}}(P_{d}h_{\mathbb{S}}[e_{d}^{-1}])\times{\mathsf{Mod}}_{\mathbb{Q}}(P_{d}h_{\mathbb{S}}[(1-e_{d})^{-1}]),$
see e.g., [Mat16, May01], where ${\mathsf{Mod}}_{\mathbb{Q}}(-)$ is shorthand notation for the $\infty$-category of modules in the rational functor category ${\mathsf{Fun}}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}$.
@@
[…]
since (leaving inclusions implicit) the left horizontal maps are the counit of the adjunction between $d$-homogeneous functors and $d$-excisive functors, while the right horizontal maps are the unit of the adjunction between $d$-excisive functors and $(d-1)$-excisive functors. Hence, any idempotent $e$ of $P_{d}h_{\mathbb{S}}$ acts by $D_{d}(e)$ on the $d$-homogeneous layer and $P_{d-1}(e)$ on the $(d-1)$-excisive approximation. The fibre sequence (1) may be identified with the fibre sequence
@@
since $P_{d-1}P_{d}\simeq P_{d-1}$ and $D_{d}P_{d}\simeq D_{d}$. Through the equivalence between (1) and (2) an idempotent $e$ of $P_{d}h_{\mathbb{S}}$ acts by $D_{d}(e)$ on the $d$-homogeneous layer and $P_{d-1}(e)$ on the $(d-1)$-excisive approximation in (2).
@@
$P_{d-1}:\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})}(P_{d}h_{\mathbb{S}},P_{d}h_{\mathbb{S}})\longrightarrow\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})}(P_{d-1}h_{\mathbb{S}},P_{d-1}h_{\mathbb{S}}).$
@@
$\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})}(P_{d-1}h_{\mathbb{S}},P_{d-1}h_{\mathbb{S}})\simeq\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{d-1}({{\mathsf{Sp}}},{{\mathsf{Sp}}})}(P_{d-1}h_{\mathbb{S}},P_{d-1}h_{\mathbb{S}}),$
@@
which sends $e_{d}$ to zero. The result follows from the observation that the action of an idempotent is determined by its action on the homotopy category. ∎
@@
$\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})}(P_{d}h_{\mathbb{S}},P_{d}h_{\mathbb{S}})\xrightarrow{\ D_{d}\ }\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})}(D_{d}h_{\mathbb{S}},D_{d}h_{\mathbb{S}}),$
@@
$\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})}(D_{d}h_{\mathbb{S}},D_{d}h_{\mathbb{S}})\simeq\operatorname{\mathsf{Hom}}_{\mathsf{Homog}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})}(D_{d}h_{\mathbb{S}},D_{d}h_{\mathbb{S}})\simeq\operatorname{\mathsf{Hom}}_{{{\mathsf{Sp}}}}(\partial_{d}h_{\mathbb{S}},\partial_{d}h_{\mathbb{S}}),$
@@
[…]
@@
[…]
@@
By 6.1 we know the last term is zero, so the first map is an equivalence. By 6.2 we know that $D_{d}h_{\mathbb{S}}[e_{d}^{-1}]\simeq D_{d}h_{\mathbb{S}}$.
@@
$\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq\prod_{1\leq i\leq d}\mathsf{Homog}_{i}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}},$
@@
By 6.3 the idempotent splitting of the $\infty$-category of $d$-excisive functors is given by
$\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}$ $\simeq{\mathsf{Mod}}_{\mathbb{Q}}(P_{d}h_{\mathbb{S}}[e_{d}^{-1}])\times{\mathsf{Mod}}_{\mathbb{Q}}(P_{d}h_{\mathbb{S}}[(1-e_{d})^{-1}])$

### v2

@@ 4. Excisive functors
A functor $F:{{\mathsf{Sp}}}^{\omega}\to{{\mathsf{Sp}}}$ is $d$-excisive if it sends strongly cocartesian $(d+1)$-cubes to cartesian cubes. One should view this as a weak form of excision, i.e., a $1$-excisive functor sends pushout diagrams to pullback diagrams. We will say that $F$ is reduced if $F(\ast)\simeq\ast$. Denote by $\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})$ the $\infty$-category of reduced $d$-excisive endofunctors of spectra [^2] For technical reasons, we must restrict the domain to the full subcategory of compact objects but we shall not belabour this point. We will also only consider reduced functors, hence drop the reduced terminology and refer to the objects of $\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})$ as $d$-excisive functors. The reduced assumption is not limiting since any functor $F:{{\mathsf{Sp}}}\to{{\mathsf{Sp}}}$ is “reduced up to a constant”, i.e., decomposes as $F(X)\simeq\overline{F}(X)\times F(\ast)$, where $\overline{F}$ is reduced..
To any $F:{{\mathsf{Sp}}}^{\omega}\to{{\mathsf{Sp}}}$, Goodwillie calculus assigns a sequence of approximations
${F}{\cdots}$$P_{d}F$$\cdots$$P_{1}F$$P_{0}F=F(\ast)$}}
in which $P_{d}F$ is the universal $d$-excisive approximation to $F$. A comprehensive account of the homotopy theory of excisive functors of spectra is given in [ABHS25, §2]. The key technical input about the $\infty$-category of $d$-excisive functors is that the category of $d$-excisive functors is closed under Day convolution of functors.
Theorem 4.1 ([ABHS25, Theorem 2.38]).
The $\infty$-category $\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})$ of $d$-excisive functors is a presentably symmetric monoidal stable $\infty$-category with monoidal structure given by Day convolution and monoidal unit $P_{d}h_{\mathbb{S}}=P_{d}\Sigma^{\infty}\operatorname{\mathsf{Map}}(\mathbb{S},-)\simeq P_{d}\Sigma^{\infty}\Omega^{\infty}$.
@@
The $\infty$-category $\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}$ of rational $d$-excisive functors is a presentably symmetric monoidal stable $\infty$-category with monoidal structure given by Day convolution and monoidal unit $L_{\mathbb{Q}}P_{d}h_{\mathbb{S}}$.
A functor $F$ is $d$-homogeneous if it is $d$-excisive and $P_{d-1}F\simeq\ast$. We will denote the $\infty$-category of $d$-homogeneous functors by $\mathsf{Homog}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})$. Since the $\infty$-category of $d$-excisive functors is closed under fibres, and a (d1)(d-1)-excisive functor is $d$-excisive, the $d$-th layer of the tower,
@@ 5. The rational Goodwillie-Burnside ring
Our attention now turns to the role of the abstract Burnside ring $A_{\mathbb{Q}}(d)$ in Goodwillie calculus, a comprehensive account of which is given in [ABHS25, §8-9]. We work rationally, as this allows us to simplify and streamline certain arguments, but integral versions of many of the statements in this section may be found in loc. cit..
@@
$\partial_{\ast}:\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})\longrightarrow{\mathsf{SymSeq}},$
@@
$\partial_{\ast}:\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})}(P_{d}h_{\mathbb{S}},P_{d}h_{\mathbb{S}})\longrightarrow\prod_{1\leq i\leq d}\operatorname{\mathsf{Hom}}_{{{\mathsf{Sp}}}}(\partial_{i}(P_{d}h_{\mathbb{S}}),\partial_{i}(P_{d}h_{\mathbb{S}})),$
@@
$\theta_{d}:\pi_{0}{\mathsf{End}}_{\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})}(P_{d}h_{\mathbb{S}})\otimes\mathbb{Q}\longrightarrow\prod_{1\leq i\leq d}\pi_{0}\operatorname{\mathsf{Hom}}_{{{\mathsf{Sp}}}}(\partial_{i}(P_{d}h_{\mathbb{S}}),\partial_{i}(P_{d}h_{\mathbb{S}}))\otimes\mathbb{Q}\cong\mathbb{Q}^{d}$
@@
between the endomorphism ring of the monoidal unit in $\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}$ and the rational Goodwillie-Burnside ring.
@@
in which the right-most map is an isomorphism by ˜2.4, hence it suffices to show that $\theta_{d}:R_{\mathbb{Q}}(d)\to\mathbb{Q}^{d}$ is an isomorphism.
By an elementary degree computation [ABHS25, Lemma 9.20], there are elements $\lambda_{m}]\in R_{\mathbb{Q}}(d)$, $1\leq m\leq d$, such that the $i$th component of $\theta_{d}[\lambda_{m}]$ is $|{\mathsf{surj}}(i,m)|$, see ˜3.5. The same example shows that the vectors $\theta_{d}[\lambda_{1}],\dots,\theta_{d}[\lambda_{d}]$ are linearly independent, hence $\theta_{d}$ is surjective for all $d$.
@@
$\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{\ast}^{1}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})}(P_{1}h_{\mathbb{S}},P_{1}h_{\mathbb{S}})\xrightarrow{\partial_{1}}\operatorname{\mathsf{Hom}}_{{{\mathsf{Sp}}}}(\partial_{1}(P_{1}h_{\mathbb{S}}),\partial_{1}(P_{1}h_{\mathbb{S}})),\simeq\operatorname{\mathsf{Hom}}_{{{\mathsf{Sp}}}}(\mathbb{S},\mathbb{S}),$
@@
For the induction step, assume that $\theta_{d-1}$ is an isomorphism. Since $\partial_{i}P_{d-1}\simeq\partial_{i}$ for $1\leq i\leq d-1$ and $\partial_{d}P_{d-1}\simeq 0$, there is a commutative diagram
$R_{\mathbb{Q}}(d)$$\mathbb{Q}^{d}$$R_{\mathbb{Q}}(d-1)$$\mathbb{Q}^{d-1}$$\scriptstyle{\theta_{d}}$$\scriptstyle{\pi_{0}(P_{d-1})\otimes\mathbb{Q}}$$\scriptstyle{(i_{d-1})^{\ast}}$$\scriptstyle{\theta_{d-1}}$
where $(i_{d-1})^{\ast}:\mathbb{Q}^{d}\to\mathbb{Q}^{d-1}$ is projection onto the first (d1)(d-1) factors. The above square induces a map
@@
The element $\lambda_{d}]$ is in the kernel of $\pi_{0}(P_{d-1})\otimes\mathbb{Q}$, hence $(\theta_{d-1}\circ\pi_{0}(P_{d-1})\otimes\mathbb{Q})[\lambda_{d}]=0$, and so, by commutativity of the above square, $\theta_{d}[\lambda_{d}]$ is in the kernel of $(i_{d-1})^{\ast}$. Since $\theta_{d}[\lambda_{d}]=(0,\dots,0,d!)$ the map $\bar{\theta}_{d}$ is non-zero and it is therefore surjective as the projection onto the first (d1)(d-1) factors, $(i_{d-1})^{\ast}:\mathbb{Q}^{d}\to\mathbb{Q}^{d-1}$ has $\ker((i_{d-1})^{\ast})\cong\mathbb{Q}$.
@@
$R_{\mathbb{Q}}(d)$$\mathbb{Q}^{d}$$A_{\mathbb{Q}}(d)$$R_{\mathbb{Q}}(d-1)$$\mathbb{Q}^{d-1}$$A_{\mathbb{Q}}(d-1)$$\scriptstyle{\theta_{d}}$$\scriptstyle{\pi_{0}(P_{d-1})\otimes\mathbb{Q}}$$\scriptstyle{(i_{d-1})^{\ast}}$$\scriptstyle{\phi_{d}}$$\scriptstyle{(i_{d-1})^{\ast}}$$\scriptstyle{\theta_{d-1}}$$\scriptstyle{\phi_{d-1}}$
in which each horizontal map is an isomorphism and the right-hand map is discussed in ˜3.6. In particular, the kernel of $\pi_{0}(P_{d-1})\otimes\mathbb{Q}$ is generated by the top idempotent in the rational Goodwillie-Burnside ring.
@@
We have already seen that the left square commutes, and the right square commutes, by definition of the map on the level of abstract Burnside rings. The claim about the rational kernel follows from combining ˜2.8 with ˜5.1 to see that the induced maps on the level of vertical kernels in the above diagram are all isomorphisms and the observation of ˜3.5 that the idempotent $e_{d}$ is a rational multiple of $[d]$. ∎
@@ 6. Splitting rational excisive functors
The Goodwillie-Burnside ring $A_{\mathbb{Q}}({\mathsf{Epi}}_{\leq d})$ associated to the category ${\mathsf{Epi}}_{\leq d}$ is equivalent to the endomorphism ring of the unit (˜5.1), so the splitting of the Goodwillie-Burnside ring (˜2.4) will induce the required splitting of the category of rational $d$-excisive functors (˜6.4).
The idempotents in the Goodwillie-Burnside ring lift to idempotents in the $\infty$-category of $d$-excisive functors by applying [Lur17, Lemma 1.2.4.6, Warning 1.2.4.8] and observing that the idempotents in the Goodwillie-Burnside ring are also idempotents in the homotopy category of $d$-excisive functors (˜5.2).
@@
by localizing with respect to the action of the idempotent $e$ and the orthogonal idempotent 1e1-e. Since the $\infty$-category of $d$-excisive functors is equivalent to modules over the monoidal unit [ABHS25, Proposition 2.35], the orthogonal idempotent splitting corresponding to $e_{d}$ may be equivalently phrased as
$\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq{\mathsf{Mod}}_{\mathbb{Q}}(P_{d}h_{\mathbb{S}})\simeq{\mathsf{Mod}}_{\mathbb{Q}}(P_{d}h_{\mathbb{S}}[e_{d}^{-1}])\times{\mathsf{Mod}}_{\mathbb{Q}}(P_{d}h_{\mathbb{S}}[(1-e_{d})^{-1}]),$
see e.g., [Mat16, May01], where ${\mathsf{Mod}}_{\mathbb{Q}}(-)$ is shorthand notation for the $\infty$-category of modules in the rational functor category ${\mathsf{Fun}}_{\ast}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}$.
@@
$D_{d}(P_{d}h_{\mathbb{S}})$$P_{d}h_{\mathbb{S}}$$P_{d-1}(P_{d}h_{\mathbb{S}})$$D_{d}(P_{d}h_{\mathbb{S}})$$P_{d}h_{\mathbb{S}}$$P_{d-1}(P_{d}h_{\mathbb{S}})$$\scriptstyle{D_{d}(e)}$$\scriptstyle{e}$$\scriptstyle{P_{d-1}(e)}$
since (leaving inclusions implicit) the left horizontal maps are the counit of the adjunction between $d$-homogeneous functors and $d$-excisive functors, while the right horizontal maps are the unit of the adjunction between $d$-excisive functors and (d1)(d-1)-excisive functors. Hence, any idempotent $e$ of $P_{d}h_{\mathbb{S}}$ acts by $D_{d}(e)$ on the $d$-homogeneous layer and $P_{d-1}(e)$ on the (d1)(d-1)-excisive approximation. The fibre sequence (1) may be identified with the fibre sequence
@@
since $P_{d-1}P_{d}\simeq P_{d-1}$ and $D_{d}P_{d}\simeq D_{d}$. Through the equivalence between (1) and (2) an idempotent $e$ of $P_{d}h_{\mathbb{S}}$ acts by $D_{d}(e)$ on the $d$-homogeneous layer and $P_{d-1}(e)$ on the (d1)(d-1)-excisive approximation in (2).
@@
$P_{d-1}:\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})}(P_{d}h_{\mathbb{S}},P_{d}h_{\mathbb{S}})\longrightarrow\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})}(P_{d-1}h_{\mathbb{S}},P_{d-1}h_{\mathbb{S}}).$
@@
$\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})}(P_{d-1}h_{\mathbb{S}},P_{d-1}h_{\mathbb{S}})\simeq\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{\ast}^{d-1}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})}(P_{d-1}h_{\mathbb{S}},P_{d-1}h_{\mathbb{S}}),$
@@
which sends $e_{d}$ to zero. The result follows from the observation that the action of an idempotent is determined by its action on the homotopy category, [Lur17, Lemma 1.2.4.6, Warning 1.2.4.8]. ∎
@@
$\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})}(P_{d}h_{\mathbb{S}},P_{d}h_{\mathbb{S}})\xrightarrow{\ D_{d}\ }\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})}(D_{d}h_{\mathbb{S}},D_{d}h_{\mathbb{S}}),$
@@
$\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})}(D_{d}h_{\mathbb{S}},D_{d}h_{\mathbb{S}})\simeq\operatorname{\mathsf{Hom}}_{\mathsf{Homog}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})}(D_{d}h_{\mathbb{S}},D_{d}h_{\mathbb{S}})\simeq\operatorname{\mathsf{Hom}}_{{{\mathsf{Sp}}}}(\partial_{d}h_{\mathbb{S}},\partial_{d}h_{\mathbb{S}}),$
@@
$\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})}(P_{d}h_{\mathbb{S}},P_{d}h_{\mathbb{S}})$$\operatorname{\mathsf{Hom}}_{\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})}(D_{d}h_{\mathbb{S}},D_{d}h_{\mathbb{S}})$$\operatorname{\mathsf{Hom}}_{{{\mathsf{Sp}}}}(\partial_{d}h_{\mathbb{S}},\partial_{d}h_{\mathbb{S}})$$\operatorname{\mathsf{Hom}}_{{{\mathsf{Sp}}}}(\partial_{d}h_{\mathbb{S}},\partial_{d}h_{\mathbb{S}})$$\scriptstyle{D_{d}}$$\scriptstyle{\partial_{d}}\scriptstyle{\simeq}$
@@
$A_{\mathbb{Q}}(d)$$\pi_{0}{\mathsf{End}}(D_{d}h_{\mathbb{S}})\otimes\mathbb{Q}$$\mathbb{Q}$$\mathbb{Q}$$\scriptstyle{\pi_{0}(D_{d})}$$\scriptstyle{\pi_{0}(\partial_{d})}\scriptstyle{\simeq}$
@@
By ˜6.1 we know the last term is zero, so the first map is an equivalence. By ˜6.2 we know that $D_{d}h_{\mathbb{S}}[e_{d}^{-1}]\simeq D_{d}h_{\mathbb{S}}$.
@@
$\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq\prod_{1\leq i\leq d}\mathsf{Homog}^{i}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}},$
@@
By ˜6.3 the idempotent splitting of the $\infty$-category of $d$-excisive functors is given by
$\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}$ $\simeq{\mathsf{Mod}}_{\mathbb{Q}}(P_{d}h_{\mathbb{S}}[e_{d}^{-1}])\times{\mathsf{Mod}}_{\mathbb{Q}}(P_{d}h_{\mathbb{S}}[(1-e_{d})^{-1}])$